OpenStudy (anonymous):

Please Help! A circle is circumscribed about a hexagon. Determine the area of the hexagon if the area outside the hexagon but inside the circle is 15 sq. cm.

OpenStudy (ranga):



OpenStudy (ranga):

Let radius of circle = r ; Area of circle = (pi)r^2 You can find the area of the triangle OAB in terms of r. Area of OAB = 1/2 * AB * OC = BC * OC = rcos(30) * rsin(30) = r^2*sqrt(3)/4 Area of hexagon = 6 * area of OAB = 6 * r^2*sqrt(3)/4 = 3/2 * sqrt(3) * r^2 Difference between the areas of hexagon and circle = (pi)r^2 - 3/2 * sqrt(3) * r^2 = 15 solve for r. put it in the area of hexagon and find its area.

OpenStudy (anonymous):

??????????????????????????????????

OpenStudy (anonymous):

um, can you use the equation button below sir? I can't really understand the flow of the solution.

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

If the math part is not clear you can still follow the description part and find the areas.

OpenStudy (ranga):

the ^ symbol means the next character is an exponent.

OpenStudy (anonymous):

[Area of OAB = 1/2 * AB * OC = BC * OC ] this part sir, how did [1/2 * AB * OC] became [BC * OC]?

OpenStudy (ranga):

* means multiplication. 1/2 * AB = BC (BC is half of AB)

OpenStudy (anonymous):

so the scope of 1/2 is only the AB excluding the OC? another question sir if I may, what about [rcos(30) * rsin(30) = r^2*sqrt(3)/4]?

OpenStudy (anonymous):

it's done by multiplying the coordinates of sin and cos of 30?

OpenStudy (anonymous):

how did you end up in rsin(30) and rcos(30)?

OpenStudy (ranga):

For example, if you have 1/2 * 6 * 8 you can associate the 1/2 with 6 or 8 and you will get the same answer. 1/2 * 6 * 8 = 3 * 8 = 24 or 1/2 * 6 * 8 = 6 * 4 = 24

OpenStudy (anonymous):

is it "SOH CAH TOA"? sir, i really apologize if I'm asking too much.

OpenStudy (anonymous):

ok sir I get the 1/2 part now.

OpenStudy (ranga):

Look up trig tables. sin(30) = 1/2 ; cos(30) = sqrt(3) / 2

OpenStudy (ranga):

sine = opposite / hypotenuse ; cosine = adjacent / hypotenuse

OpenStudy (anonymous):

yes sir I know that trigo part but the transition in [BC * OC = rcos(30) * rsin(30)] is what I'm confused at.

OpenStudy (ranga):

sin(30) = BC / OB = BC / r Therefore, BC = r * sin(30)

OpenStudy (ranga):

Angle BOC = 30 degrees (half of 60 degrees)

OpenStudy (anonymous):

aaaah. I understand it now sir :D

OpenStudy (ranga):

alright.

OpenStudy (anonymous):

now all I need is to solve for r and get the area of the hexagon.

OpenStudy (ranga):

yes. But I will be logging off now but hopefully someone will be here to help if needed.

OpenStudy (anonymous):

Thank you sir. @ranga

OpenStudy (ranga):

you are welcome.

OpenStudy (anonymous):

I can't solve for r. too complex.

OpenStudy (anonymous):

The area of region between the circle and the hexagon is given by ..Area of the circle(pir^2)-(area of the hexagon(1/2 r^2sin60x6 (because hexagon forms six identical triangles) there, 15=pi(3.142) x r^2 -(0.5 x r^2sin60 x6) 15=3.142r^2 -2.598r^2 15=0.544r^2 15/0.544 =r^2 27.58=r^2 square root both sides to get radius Please verify your email address. We've sent you an email with a verification link inside. OpenStudy Home Invite omarani1336 Settings Log out SmartScore ← 96 members online 0 replying 2 viewing omarani1336 I am a little bit confused of how to obtain ratios from the unit cycle plz help a few moments ago Type your reply You have an open question: I am a little bit confused of how to obtain ratios from the unit cycle plz help Asked a few moments ago No Medals Yet lowcard2: help?!!!!! 3 viewing No Medals Yet ECEstudent9405: Please Help! A circle is circumscribed about a hexagon. Determine the area of the hexagon if the area outside the hexagon but inside the ci… 1 replying right now... 1 viewing No Medals Yet Ratedover: Hey! I know this is math but I've been trying English but I have't had any luck with it. Could anyone give me some feedback on my essay? It … 2 viewing No Medals Yet stupidinmath: Two regular pentagons are constructed one on the hypotenuse and one on a leg of an isosceles right triangle. What is the ratio of the areas … 1 viewing No Medals Yet Geometry omarani1336: I am a little bit confused of how to obtain ratios from the unit cycle plz help 2 viewing No Medals Yet Abbles: Algebra homework help? updated a few moments ago No Medals Yet brendafn: change to exponential form y=log base(6) X 2 viewing 2 Medals Yttrium: Stats question: I just wanna verify if I got the correct answer. "Three coins, consisting two 10$, three 5$, and five 1$, are to be selecte… 1 viewing No Medals Yet stupidinmath: Two regular pentagons are constructed one on the hypotenuse and one on a leg of an isosceles right triangle. What is the ratio of the areas … 1 viewing 1 Medal NotTim: Problem 6. 6. (2 pts) Construct a 98% confidence interval for the population slope b. b = 34, Se = 7; SSx = 45;n = 11 … 1 viewing No Medals Yet No more questions. Working hard or hardly working? Privacy Policy Terms and Conditions Code of Conduct Mathematics Search Subjects 95 Online 58 Online 53 Online 47 Online 46 Online 46 Online 41 Online 39 Online 39 Online 27 Online Loading more… Create a New Subject Enter Subject Name

OpenStudy (anonymous):

then when you get the radius you can simply find the area of the hexagon by using the formula 1/2 r^2sin60 x6 0.5x(5.3)^2sin60 x 6=72.98 sq cm

OpenStudy (anonymous):

hey are you getting my response/